Hearing the siren of an approaching fire truck, you pull over to the side of the road and stop. As the truck approaches, you hear a tone of 470 Hz ; as the truck recedes, you hear a tone of 410 Hz .

How much time will it take for the truck to get from your position to the fire 6.0 km away, assuming it maintains a constant speed?
Answer #1

I am assuming speed of sound to be 340 m/s.

equate the observed frequencies to the frequency of the source

freceding ( v + vs) / v = fapproach ( v - vs) / v

freceding ( v + vs) = fapproach ( v - vs)

freceding * v + freceding * vs =  fapproach * v - fapproach *vs

Solving for vs , we get

vs = v(fapproach - freceding ) / (fapproach + freceding )

vs = 340 (470 - 410) / 470 + 410

vs = 23.18 m/s

-----------------------------------------------------------------------------

Now, we know the speed, and we have the distance

so,

time taken = distance / speed

time = 258.8 seconds or 4.31 minutes